On a test, the number of questions Tyler answered correctly was 4 more than twice the number that he answered incorrectly. If the ratio of correct answers to incorrect answers is 8:3 how many questions did he get right?
Please show work and explain.

Answer by TimothyLamb(4379)

---
r/w = 8/3
r/w = (2x + 4)/x
---
(2x + 4)/x = 8/3
2x + 4 = (8/3)x
4 = (8/3)x - 2x
4 = (8/3)x - (6/3)x
4 = (2/3)x
x = 4(3/2)
---
answer:
x = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php